Mathematics 243 – Algebraic Structures Definitions Quiz 1 – September 8, 2006

Your Name:

A) (5) Let f : A → B be a mapping. What is the definition of the property that f is injective (one-to-one).

B) (5) Let A = {−4, −2, 0, 2, 4, 6, 8} and B = {x ∈ Z | 1 ≤ x ≤ 5}. What is the set A ∩ B?